2. There are two points in a constant electric field with a magnitude 80 V/m which is directed parallel to the x-axis. Point l is at (xi,y, ) = (3, 4) in m, and point 2 is at (x2,y2)-(12,9) in m. The potential at point 1 is 1000 V. Calculate the potential at point 2. Calculate the work required to move a negative charge ofq562 uC from point 1 to point 2.
Answer #1

electric field = 80 V/m in x-direction

seperation between the points = sqrt( ( 12-3)^2 + (9-4)^2)

= 10.3 m

V2 - V1 = E*d

V2 - 1000 = 80*10.3

V2 = 1823.65 V

work done = Charge * potential difference

= -562*10^-6 * 1823.65

= -1.02 J